The following is simplified aggregate planning information (no hiring or layoff is expected):

July

August

September

Beginning inventory, units

105

Demand forecast, units

610

530

720

Safety stock, units

Production requirements, units

Workers required

20

20

20

New workers hired

0

0

0

Workers laid off

0

0

0

Actual production, units

650

650

650

Ending inventory, units

Consider that these months have the same number of workdays. There will be no new hiring or layoffs during the 3-month period. Also, consider that safety stock is equal to 10% of each monthly demand.

a. What aggregate planning strategy is being used in this case? When compare with chase strategy, please explain the pros and cons of each strategy.

b. For each month, calculate safety stock, production requirements, beginning inventory (except for July), and ending inventory. Show your calculations, and input results in the table above.

c. Calculate the total cost of inventory for the entire three-month plan. The cost is $2 per month to hold one unit in inventory.
Answer #1

a.The aggregate planning strategy used in this case is level strategy because the production level is constant and is equal to 650 units.

The pros of this strategy:

  • The output is stable and also the workforce is held constant which enhances the employee morale.
  • The total cost is lower when compared to chase planning strategy.

The cons of this strategy :

  • Higher inventory costs
  • Higher labor utilization.

b.Safety stock =10 % of demand

Production requirement =Demand forecast+Safety stock

Production requirement for July=B3+B4=671

(Drag this cell formula to the right to the rest of the months)

Ending inventory for July=B2+B9-B3+B4=206

(Drag this cell formula to the right to the rest of the months)

Beginning inventory of August=Ending inventory of July and so on.

B13 - fic -SUM(B11:011) A 1 July 105 August September 206 379 610 530 720 53. 583 792 20 72 2 Beginning Inventory | 3 Demand

c.Inventory cost for July=B10*2=412

(Drag this cell formula to the right to the rest of the months)

Total inventory cost=SUM(B11:D11)=$ 1932